OpenStudy (anonymous):

WILL GIVE MEDAL!!! find the length of the unknown side of the right triangle with sides a,b, and c. where c is the hypotenuse. a=√9, c=9

OpenStudy (mertsj):

a^2+b^2=c^2 Plug into that and find b. b is the missing side.

OpenStudy (anonymous):

use pyth.theorem...base sqr.+altd sqr=hyp sqr

OpenStudy (anonymous):

(root 9)^2+b^2=9^2............9+b^2=81..........b^2=81-9=72

OpenStudy (chmvijay):

still finding difficulty to this problem @lashell123